OpenStudy (anonymous):

How many different three-letter arrangements can be formed using the letters in the word ABSOLUTE if each letter is used only once? A. 56 B. 112 C. 168 D. 336

OpenStudy (anonymous):

Woo!! last question!! ^^

OpenStudy (watchmath):

There are 8 letters and we want to choose 3 among them. So 8 choose 3, i.e \(\frac{8!}{5!3!}=\frac{6\times 7\times 8}{1\times 2\times 3}=56\)

OpenStudy (anonymous):

simplify the expression: (√7+√385)(√55+√11)

OpenStudy (anonymous):

Ohhhh okie dokie (: Thanks a but load